### v2

@@ Abstract
Large Language Models (LLMs), such as the GPT-4 and LLaMA families, have demonstrated considerable success across diverse tasks, including multiple-choice questions (MCQs). However, these models exhibit a positional bias, particularly an even worse “anchored bias” in the GPT-2 family, where they consistently favour the first choice ‘A’ in MCQs during inference. This anchored bias challenges the integrity of GPT-2’s decision-making process, as it skews performance based on the position rather than the content of the choices in MCQs. In this study, we utilise the mechanistic interpretability approach to identify the internal modules within GPT-2 models responsible for this bias. We focus on the Multi-Layer Perceptron (MLP) layers and attention heads, using the “logit lens” method to trace and modify the specific value vectors that contribute to the bias. By updating these vectors within MLP and recalibrating attention patterns to neutralise the preference for the first choice ‘A’, we effectively mitigate the anchored bias. Our interventions not only mitigate the bias but also improve the overall MCQ prediction accuracy for the GPT-2 family across various datasets. This work represents the first comprehensive mechanistic analysis of anchored bias in MCQs within the GPT-2 models, introducing targeted, minimal-intervention strategies that significantly enhance GPT2 model robustness and accuracy in MCQs. Our code is available at https://github.com/ruizheliUOA/Anchored_Bias_GPT2.
@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable capabilities across a wide array of tasks, including multiple-choice question (MCQ) (Robinson and Wingate,, 2022), which are largely attributed to the advancements in the Transformer backbone. These models not only excel at reasoning but also demonstrate significant inductive capabilities, which make them highly effective in different domains (Chen et al.,, 2023; Team et al.,, 2023; Anil et al.,, 2023).
Despite their success, recent studies have uncovered a notable flaw: LLMs exhibit a positional bias when tasked with MCQs. Specifically, the performance of these models (e.g., LLaMA (Touvron et al., 2023a, ), LLaMA2 (Touvron et al., 2023b, ), GPT-4 (Achiam et al.,, 2023)) varies significantly depending on the position of the correct answer within the given choices (Pezeshkpour and Hruschka,, 2023; Zheng et al., 2024a, ). We further observe that this vulnerability to positional bias is even worse in the GPT-2 family, ranging from GPT2-Small-124M to GPT2-XL-1.5B (Radford et al.,, 2019). Our investigations reveal that GPT-2 models consistently favour the first choice ‘A’, regardless of the actual position in the input MCQ prompt where the correct answer choice is placed, which we term as “anchored bias” in Fig. 1.
Previous work primarily mitigated positional bias in MCQ by analysing the impact of different prompt structures (Pezeshkpour and Hruschka,, 2023) or by estimating different datasets’ prior bias based on test samples (Zheng et al., 2024a, ). Such approaches often remain superficial, merely altering the prompt presentation, or lacking a comprehensive analysis of fundamental reasons. Remarkably, there has been a lack of investigation into the internal mechanisms of LLMs that contribute to the anchored bias and strategies to mitigate it without the need for prompt engineering or prior estimation.
@@
We apply mechanistic interpretability to reverse-engineer the internal workings of the GPT-2 family to understand the origins and extent of the anchored bias. We quantitatively demonstrate that the GPT-2 Small, Medium, Large, and XL models exhibit this anchored bias with significant regularity across various MCQ datasets, ranging from 2 choices to 5 choices settings. Our detailed analysis using the “logit lens” (Nostalgebraist,, 2020) approach localises Multi-Layer Perceptron (MLP) layers with specific dimensionality and attention heads that disproportionately influence this anchored bias. We find that certain value vectors in the MLP, which inherently harbour this bias, and specific attention heads pay more weight on the ‘A’ position over the correct answer choice positions in the input prompt.
@@
Inspired from (Geva et al.,, 2021, 2022) where MLPs can be treated as key-value memories, we use a straightforward yet potent method (Dai et al.,, 2022) to update these critical value vectors in the MLP, effectively mitigate the anchored bias. This adjustment not only mitigates the anchored bias but also enhances the overall MCQ prediction accuracy over 70% averaged across various MCQ datasets and all GPT-2 family. Additionally, we propose a novel strategy to recalibrate the attention patterns by swapping the attention weight between the anchored position and the correct answer choice position. This strategy also mitigates the anchored bias to a certain extent, especially for the classification accuracy improvement of the Indirect Obejct Identification (IOI) dataset (Wang et al.,, 2022) over 90% on GPT2-Medium. Finally, we trace the full anchored bias circuit of each GPT2 model, which includes all attention heads and MLPs contributing to this bias.
@@
In conclusion, to our best knowledge, this work is the first comprehensive mechanistic analysis of the intrinsic anchored bias in MCQ tasks across the entire GPT-2 family. By identifying and rectifying the critical value vectors within MLP and attention heads responsible for this bias, we introduce novel, minimal-intervention strategies that significantly reduce GPT-2 models’ vulnerability and enhance robustness against anchored bias in MCQ task.
@@ 2 Background: Large Language Models and Mechanistic Interpretability
Figure 2: Left: the MLP and attention modules of LLMs, where the input prompt is encoded via $W_{E}$, and then the processed information via attention and MLP layer is accumulated back to the residual stream $\mathbf{X}^{\ell}$ at layer $\ell$. Finally, the residual stream at $L$ layer is unembedded as logits and normalised as a probability distribution for next token prediction. Right: logit lens (Nostalgebraist,, 2020) is used to investigate the contribution of attention pattern and MLP module for the next token prediction.
@@
We focus on the autoregressive Transformer-based LLM architecture (Vaswani et al.,, 2017) based on prior works (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022; Meng et al., 2022a, ; Meng et al., 2022b, ; Yuksekgonul et al.,, 2024) with simplifications in certain explanations. Given an input prompt containing $T$ tokens $(t_{1},\ldots,t_{T})$ and each token $t_{i}$ belonging to a vocabulary $\mathcal{V}$, tokens are initially encoded by $d$-dimensional vectors $\mathbf{x}_{i}^{0}\in\mathbb{R}^{d}$ using an embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
As shown in Fig. 2, the architecture has $L$ layers, and each layer consists of attention and MLP modules, which transform token embeddings to residual streams $(\mathbf{x}_{1}^{\ell},\ldots,\mathbf{x}_{T}^{\ell})\in\mathbf{X}^{\ell}$ at layer $\ell$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$. The residual stream at layer $\ell$ is a place where all attention and MLP modules at layer $\ell$ read from and write to (Elhage et al.,, 2021), and it is updated by the following equation for token $i$ at layer $\ell$:
@@
The attention module mainly updates each token residual stream $\mathbf{x}_{i}^{l-1}$ by attending to all previous tokens in parallel. Specifically, the attention module contains $QK$ and $OV$ circuits, where the former operates $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ matrices and the latter operates $W_{O},W_{V}\in\mathbb{R}^{d\times d}$ matrices, respectively. Normally, $QK$ circuit determines the attention pattern $A^{\ell}$, i.e., where information is moved to and from the residual stream. $OV$ circuit further determines the attention output $\mathbf{a}_{i}^{\ell}$ based on the fixed attention pattern, i.e., what information is from the previous tokens’ position to the current token position (Elhage et al.,, 2021):
@@
where $\mathbf{a}_{i,j}^{\ell}$ indicates the attention contribution from token $i$ to token $j$, and $\mathbf{a}_{i}^{\ell}=\sum_{j=1}^{T}\mathbf{a}_{i,j}^{\ell}$. $\mathbf{r}_{i,j}^{l,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$, and $\mathbf{r}_{i}^{\ell,h}=\sum_{j=1}^{T}\mathbf{r}_{i,j}^{\ell,h}$ (See Appendix C for detailed explanations about attention module).
MLP module is normally treated as key-value memories (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022), where columns of $W_{\text{in}[:,i]}^{\ell}$ and rows of $W_{\text{out}[i,:]}^{\ell}$ are viewed as keys and values in Fig. 2, respectively. Given the input $\mathbf{x}_{i}^{\ell-1}$, the keys of MLP produce a vector of cofficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, and they weights the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$ (See Appendix C for detailed introduction about MLP):
@@
Logit lens is a mechanistic interpretability approach to investigate the contribution of the intermediate layer representation in the autoregressive Transformer-based LLMs (Nostalgebraist,, 2020). Based on the architecture of LLMs above, the $\mathcal{P}(t_{T+1}|t_{1:T})$ at layer $L$ is the production of linear softmax of logits unembedded via $W_{U}$, which is the sum of input $\mathbf{x}_{i}^{0}$ and attention and MLP contributions at each layer $\ell$. Therefore, logit lens can be used to measure the weighted attention value of each head $\mathbf{r}_{i}^{\ell,h}\in\mathbb{R}^{d}$, each weighted value vector $\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n}\in\mathbb{R}^{d}$ at $n$-th dimensionality in MLP and intermediate residual stream $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$ for token $i$:
$\text{logit}_{i}^{\ell,h}(\mathbf{r}_{i}^{\ell,h})=W_{U}\sigma(\mathbf{r}_{i}^{\ell,h})\quad\text{logit}_{i}^{\ell,n}(\mathbf{m}_{i}^{\ell,n})=W_{U}\sigma(\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n})\quad\text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})=W_{U}\sigma(\mathbf{x}_{i}^{\ell})$ (5)
@@ 3 Preliminaries: Zero-shot Learning with MCQs
To comprehensively verify and evaluate the anchored bias of GPT2 family, we consider 5 datasets, which include different numbers of choices from 2 to 5. Indirect Object Identification (IOI) Wang et al., (2022) and Greater-than task (Greater) Hanna et al., (2024): These two datasets have been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024; Hanna et al.,, 2024). However, we found that GPT2 family immediately fails these tasks if the input prompt is formatted as MCQ in Fig. 1, where the incorrect subject of the last clause or incorrect years is placed in the ‘A’ choice and the prediction is always anchored at incorrect choice ‘A’. Logical Deduction of the Big-Bench (LD) [^2] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al., (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. ARC-Challenge (ARC) Clark et al., (2018) and CommensenseQA (CSQA) Talmor et al., (2019) are commonly-used MCQ benchmarks to evaluate LLMs. For each dataset, we split into 90% Infer. set for anchored bias discovering and mitigation, and 10% Eva. set to evaluate the modified GPT2 model performance. For models, we comprehensively evaluate the GPT2 family, i.e., GPT2-Small-124M, GPT2-Medium-355M, GPT2-Large-774M, and GPT2-XL-1.5B Radford et al., (2019) (See Appendix D for detailed introduction of each dataset).
@@
We use logit lens (Nostalgebraist,, 2020) introduced in § 2 to localise the specific layer of MLP and specific attention head which contribute to the anchored bias (more details in § 4). Moreover, we use MLP contribution (Geva et al.,, 2022) to locate the specific dimensionality from $W_{\text{out}}^{\ell}$ which leads to anchored bias. Regarding mitigating anchored bias, we use classification accuracy to evaluate whether the anchored bias can be mitigated and GPT2 family can successfully predict correct choices in MCQ.

### v3

@@ Abstract
Large Language Models (LLMs), such as the GPT-4 and LLaMA families, have demonstrated considerable success across diverse tasks, including multiple-choice questions (MCQs). However, these models exhibit a positional bias, particularly an even worse “anchored bias” in the GPT-2 family, where they consistently favour the first choice ‘A’ in MCQs during inference. This anchored bias challenges the integrity of GPT-2’s decision-making process, as it skews performance based on the position rather than the content of the choices in MCQs. In this study, we utilise the mechanistic interpretability approach to identify the internal modules within GPT-2 models responsible for this bias. We focus on the Multi-Layer Perceptron (MLP) layers and attention heads, using the “logit lens” method to trace and modify the specific value vectors that contribute to the bias. By updating these vectors within MLP and recalibrating attention patterns to neutralise the preference for the first choice ‘A’, we effectively mitigate the anchored bias. Our interventions not only mitigate the bias but also improve the overall MCQ prediction accuracy for the GPT-2 family across various datasets. This work represents the first comprehensive mechanistic analysis of anchored bias from the failing cases in MCQs within the GPT-2 models, introducing targeted, minimal-intervention strategies that significantly enhance GPT2 model robustness and accuracy in MCQs. Our code is available at https://github.com/ruizheliUOA/Anchored_Bias_GPT2.
Anchored Answers: Unravelling Positional Bias
in GPT-2’s Multiple-Choice Questions
Ruizhe Li1, Yanjun Gao2 1Department of Computing Science, University of Aberdeen, UK 2University of Colorado Anschutz Medical Campus, United States Correspondence: ruizhe.li@abdn.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable capabilities across a wide array of tasks, including multiple-choice question (MCQ) (Robinson and Wingate, 2023), which are largely attributed to the advancements in the Transformer backbone. These models not only excel at reasoning but also demonstrate significant inductive capabilities, which make them highly effective in different domains (Hu et al., 2023b; Chen et al., 2023; Hu et al., 2023a; Team et al., 2023; Anil et al., 2023; Hu et al., 2024a, b; Li et al., 2025; Liu et al., 2025; Ji et al., 2024; Gao et al., 2025).
Despite their success, recent studies have uncovered a notable flaw: LLMs exhibit a positional bias when tasked with MCQs. Specifically, the performance of these models (e.g., LLaMA (Touvron et al., 2023a), LLaMA2 (Touvron et al., 2023b), GPT-4 (Achiam et al., 2023)) varies significantly depending on the position of the correct answer within the given choices (Pezeshkpour and Hruschka, 2024; Zheng et al., 2024a). We further observe that this vulnerability to positional bias is even worse in the GPT-2 family, ranging from GPT2-Small-124M to GPT2-XL-1.5B (Radford et al., 2019). Our investigations reveal that GPT-2 models consistently favour the first choice ‘A’, regardless of the actual position in the input MCQ prompt where the correct answer choice is placed, which we term as “anchored bias” in Fig. 1.
Previous work primarily mitigated positional bias in MCQ by analysing the impact of different prompt structures (Pezeshkpour and Hruschka, 2024) or by estimating different datasets’ prior bias based on test samples (Zheng et al., 2024a). Such approaches often remain superficial, merely altering the prompt presentation, or lacking a comprehensive analysis of fundamental reasons. While Lieberum et al. (2023); Wiegreffe et al. (2025) investigated positional bias-related problem using mechanistic interpretability, they mainly focus on success cases. There has been a lack of investigation into the internal mechanisms of LLMs from the failing cases that contribute to the anchored bias and strategies to mitigate it without the need for prompt engineering or prior estimation.
@@
We apply mechanistic interpretability to reverse-engineer the internal workings of the GPT-2 family to understand the origins and extent of the anchored bias in the failing cases. We quantitatively demonstrate that the GPT-2 Small, Medium, Large, and XL models exhibit this anchored bias with significant regularity across various MCQ datasets, ranging from 2-choice to 5-choice settings. Our detailed analysis using the “logit lens” (Nostalgebraist, 2020) approach localises Multi-Layer Perceptron (MLP) layers with specific dimensionality and attention heads that disproportionately influence this anchored bias. We find that certain value vectors in the MLP, which inherently harbour this bias, and specific attention heads pay more weight on the ‘A’ position over the correct answer choice positions in the input prompt.
@@
Inspired from (Geva et al., 2021, 2022) where MLPs can be treated as key-value memories, we use a straightforward yet potent method (Dai et al., 2022) to update these critical value vectors in the MLP, effectively mitigating the anchored bias. This adjustment not only mitigates the anchored bias but also enhances the overall MCQ prediction accuracy over 70% averaged across various MCQ datasets and all GPT-2 family. Additionally, we propose a novel strategy to recalibrate the attention patterns by swapping the attention weight between the anchored position and the correct answer choice position. This strategy also mitigates the anchored bias to a certain extent, especially for the classification accuracy improvement of the Indirect Object Identification (IOI) dataset (Wang et al., 2023) over 90% on GPT2-Medium. Finally, we trace the full anchored bias circuit of each GPT2 model, which includes all attention heads and MLPs contributing to this bias.
@@
In conclusion, to the best of our knowledge, this work is the first comprehensive mechanistic analysis of the intrinsic anchored bias from the failing cases in MCQ tasks across the entire GPT-2 family. By identifying and rectifying the critical value vectors within MLP and attention heads responsible for this bias, we introduce novel, minimal-intervention strategies that significantly reduce GPT-2 models’ vulnerability and enhance robustness against anchored bias in the MCQ task.
2 Related Work
Several studies have documented the effects of positional bias on LLM accuracy in MCQs. Pezeshkpour and Hruschka (2024) found a "sensitivity gap" in models like GPT-4, where positional bias can decrease performance by up to 75% in a zero-shot setting, and they improved accuracy with new calibration strategies. Wang et al. (2024) also noted the impact of option order on GPT-4’s scores, enhancing accuracy through a calibration framework including multiple evidence and balanced position adjustments, along with human involvement. Zheng et al. (2024a) addressed "selection bias" where LLMs disproportionately favour certain options, introducing a debiasing method, PriDe, that adjusts predictions during inference. Wang et al. (2025) explored performance changes from reordering answer options, confirming that this impacts understanding. Turpin et al. (2023) and Zheng et al. (2024b) explore the effects of positional bias in LLMs, showing how it skews Chain-of-Thought generation and evaluator judgments, and emphasise the need for strategies to detect and mitigate these biases. However, those studies did not analyse such bias within models and further identify which component is relevant. Recently,  Lieberum et al. (2023) analyses final-token attention heads and identifies a subset “correct letter heads”, which focus on earlier answer symbols to promote the correct choice based on its order (mainly for A/B/C/D). However, their findings are based solely on MMLU task using the closed-source Chinchilla-70B. Wiegreffe et al. (2025) investigates how successful models perform formatted MCQs with symbol binding internally rather than focusing on the failure cases when models have positional bias. Compared to (Lieberum et al., 2023; Wiegreffe et al., 2025) focusing on the success cases, our work is focused on the failing cases, and we believe that these complementary perspectives help form a more complete understanding of how LLMs handle MCQs.
@@ 3 Background: Large Language Models and Mechanistic Interpretability
Figure 2: Left: the MLP and attention modules of LLMs, where the input prompt is encoded via $W_{E}$, and then the processed information via attention and MLP layer is accumulated back to the residual stream $\mathbf{X}^{\ell}$ at layer $\ell$. Finally, the residual stream at $L$ layer is unembedded as logits and normalised as a probability distribution for next token prediction. Right: logit lens (Nostalgebraist, 2020) is used to investigate the contribution of attention pattern and MLP module for the next token prediction.
@@
We focus on the autoregressive Transformer-based LLM architecture (Vaswani et al., 2017) based on prior works (Geva et al., 2021, 2022; Elhage et al., 2021; Dai et al., 2022; Meng et al., 2022, 2023; Yuksekgonul et al., 2024) with simplifications in certain explanations. Given an input prompt containing $T$ tokens $(t_{1},\ldots,t_{T})$ and each token $t_{i}$ belonging to a vocabulary $\mathcal{V}$, tokens are initially encoded by $d$-dimensional vectors $\mathbf{x}_{i}^{0}\in\mathbb{R}^{d}$ using an embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
As shown in Fig. 2, the architecture has $L$ layers, and each layer consists of attention and MLP modules, which transform token embeddings to residual streams $(\mathbf{x}_{1}^{\ell},\ldots,\mathbf{x}_{T}^{\ell})\in\mathbf{X}^{\ell}$ at layer $\ell$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$. The residual stream at layer $\ell$ is a place where all attention and MLP modules at layer $\ell$ read from and write to (Elhage et al., 2021), and it is updated by the following equation for token $i$ at layer $\ell$:
@@
The attention module mainly updates each token residual stream $\mathbf{x}_{i}^{l-1}$ by attending to all previous tokens in parallel. Specifically, the attention module contains $QK$ and $OV$ circuits, where the former operates $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ matrices and the latter operates $W_{O},W_{V}\in\mathbb{R}^{d\times d}$ matrices, respectively. Normally, $QK$ circuit determines the attention pattern $A^{\ell}$, i.e., where information is moved to and from the residual stream. $OV$ circuit further determines the attention output $\mathbf{a}_{i}^{\ell}$ based on the fixed attention pattern, i.e., what information is from the previous tokens’ position to the current token position (Elhage et al., 2021):
@@
where $\mathbf{a}_{i,j}^{\ell}$ indicates the attention contribution from token $i$ to token $j$, and $\mathbf{a}_{i}^{\ell}=\sum_{j=1}^{T}\mathbf{a}_{i,j}^{\ell}$. $\mathbf{r}_{i,j}^{l,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$, and $\mathbf{r}_{i}^{\ell,h}=\sum_{j=1}^{T}\mathbf{r}_{i,j}^{\ell,h}$ (See Appendix B for detailed explanations about attention module).
MLP module is normally treated as key-value memories (Geva et al., 2021, 2022; Elhage et al., 2021; Dai et al., 2022), where columns of $W_{\text{in}[:,i]}^{\ell}$ and rows of $W_{\text{out}[i,:]}^{\ell}$ are viewed as keys and values in Fig. 2, respectively. Given the input $\mathbf{x}_{i}^{\ell-1}$, the keys of MLP produce a vector of cofficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, and they weights the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$ (See Appendix B for detailed introduction about MLP):
@@
Logit lens is a mechanistic interpretability approach to investigate the contribution of the intermediate layer representation in the autoregressive Transformer-based LLMs (Nostalgebraist, 2020). Based on the architecture of LLMs above, the $\mathcal{P}(t_{T+1}|t_{1:T})$ at layer $L$ is the production of linear softmax of logits unembedded via $W_{U}$, which is the sum of input $\mathbf{x}_{i}^{0}$ and attention and MLP contributions at each layer $\ell$. Therefore, logit lens can be used to measure the weighted attention value of each head $\mathbf{r}_{i}^{\ell,h}\in\mathbb{R}^{d}$, each weighted value vector $\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n}\in\mathbb{R}^{d}$ at $n$-th dimensionality in MLP and intermediate residual stream $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$ for token $i$:
$\begin{split}\text{logit}_{i}^{\ell,h}(\mathbf{r}_{i}^{\ell,h})&=W_{U}\sigma(\mathbf{r}_{i}^{\ell,h})\\ \text{logit}_{i}^{\ell,n}(\mathbf{m}_{i}^{\ell,n})&=W_{U}\sigma(\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n})\\ \text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})&=W_{U}\sigma(\mathbf{x}_{i}^{\ell})\end{split}$ (5)
@@ 4 Preliminaries: Zero-shot Learning with MCQs
To comprehensively verify and evaluate the anchored bias of GPT2 family, we consider 5 datasets, which include different numbers of choices from 2 to 5. Indirect Object Identification (IOI) Wang et al. (2023) and Greater-than task (Greater) Hanna et al. (2023): These two datasets have been verified that GPT2 family works well (Wang et al., 2023; Merullo et al., 2024; Hanna et al., 2023). However, we found that GPT2 family immediately fails these tasks if the input prompt is formatted as MCQ in Fig. 1, where the incorrect subject of the last clause or incorrect years is placed in the ‘A’ choice and the prediction is always anchored at incorrect choice ‘A’. Logical Deduction of the Big-Bench (LD) [^2] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al. (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. ARC-Challenge (ARC) Clark et al. (2018) and CommensenseQA (CSQA) Talmor et al. (2019) are commonly-used MCQ benchmarks to evaluate LLMs. For each dataset, we split into 90% Infer. set for anchored bias discovering and mitigation, and 10% Eva. set to evaluate the modified GPT2 model performance without accessing the gold labels. For models, we comprehensively evaluate the GPT2 family, i.e., GPT2-Small-124M, GPT2-Medium-355M, GPT2-Large-774M, and GPT2-XL-1.5B Radford et al. (2019) (See Appendix C for a detailed introduction of each dataset).
@@
We use logit lens (Nostalgebraist, 2020) introduced in § 3 to localise the specific layer of MLP and specific attention head which contribute to the anchored bias (more details in § 5). Moreover, we use MLP contribution (Geva et al., 2022) to locate the specific dimensionality from $W_{\text{out}}^{\ell}$ which leads to anchored bias. Regarding mitigating anchored bias, we use classification accuracy to evaluate whether the anchored bias can be mitigated and GPT2 family can successfully predict correct choices in MCQ.